find the inverse of h(x)=5x+6/4x-9

Answer by stanbon(57377)

find the inverse of h(x)=5x+6/4x-9
---
Interchange x and y:
x = (5y+6)/(4y-9)
Cross-multiply:
4xy - 9x = 5y+6
------------------------
Solve for "y"
y(4x-5) = 6-9x
y = (6-9x)/4x-5)
That is the inverse.
===========================
Cheers,
Stan H.